# Allowing humans to interactively guide machines where to look does not always improve human-AI team’s classification accuracy

### v2

@@ Abstract
Via thousands of papers in Explainable AI (XAI), attention maps [50] and feature attribution maps [9] have been established as a common means for finding how important each input feature is to an AI’s decisions. It is an interesting, unexplored question whether allowing users to edit the feature importance at test time would improve a human-AI team’s accuracy on downstream tasks. In this paper, we address this question by leveraging CHM-Corr, a state-of-the-art, ante-hoc explainable classifier [48] that first predicts patch-wise correspondences between the input and training-set images, and then base on them to make classification decisions. We build CHM-Corr++, an interactive interface for CHM-Corr, enabling users to edit the feature attribution map provided by CHM-Corr and observe updated model decisions. Via CHM-Corr++, users can gain insights into if, when, and how the model changes its outputs, improving their understanding beyond static explanations. However, our user study with 18 users who performed 1,400 decisions finds no statistical significance that our interactive approach improves user accuracy on CUB-200 bird image classification over static explanations. This challenges the hypothesis that interactivity can boost human-AI team accuracy [46, 47, 42, 44, 34, 30, 16, 32] and raises needs for future research. We open-source CHM-Corr++, an interactive tool for editing image classifier attention (see an interactive demo here). We release code and data on github.
@@ 1 Introduction
Despite much attention from the community, the practical utility of Explainable AI (XAI) tools in downstream applications (e.g., image classification [20, 24, 37, 43]) remains limited, hindering human-AI collaboration in real-world settings. A major limitation is that there is no interface for humans to provide feedback to the AI for the model to update its decisions, which could change users’ thoughts and final decisions. For example, feature attribution maps [9, 15] and example-based explanations [37] are among the most popular XAI methods that offer insights into “what a model is looking at” and which real examples support a model decision, respectively. However, they only offer a static, one-time explanation of the input. Here, we test allowing users to edit the “attention” input to the model and observe updated model decisions iteratively until they are ready to make decisions (Fig. 1).
@@
We perform our study on CHM-Corr, a recent explainable CUB-200 bird classifier [48] that combines the best of both worlds by first finding, at the patch level, how the input image is similar to the nearest training-set examples, and then using these patch-wise correspondences to predict the image label (see Fig. 2b and Supp. A3 for examples of CHM-Corr explanations). CHM-Corr explanations enable users to achieve state-of-the-art human-AI team accuracy in bird identification on CUB-200 [48].
We build an interactive interface called CHM-Corr++ for CHM-Corr, allowing users to manipulate the “attention” of the CHM-Corr classifier by selecting the set of patches that the classifier uses in decision-making (Fig. 1). By iteratively telling the model where to look, and observing the changes in the output space (see Fig. 2), users could better understand the AI model and make more accurate decisions.

### v3

@@ Abstract
Via thousands of papers in Explainable AI (XAI), attention maps [50] and feature importance maps [9] have been established as a common means for finding how important each input feature is to an AI’s decisions. It is an interesting, unexplored question whether allowing users to edit the feature importance at test time would improve a human-AI team’s accuracy on downstream tasks. In this paper, we address this question by leveraging CHM-Corr, a state-of-the-art, ante-hoc explainable classifier [48] that first predicts patch-wise correspondences between the input and training-set images, and then bases on them to make classification decisions. We build CHM-Corr++, an interactive interface for CHM-Corr, enabling users to edit the feature importance map provided by CHM-Corr and observe updated model decisions. Via CHM-Corr++, users can gain insights into if, when, and how the model changes its outputs, improving their understanding beyond static explanations. However, our study with 18 expert users who performed 1,400 decisions finds no statistical significance that our interactive approach improves user accuracy on CUB-200 bird image classification over static explanations. This challenges the hypothesis that interactivity can boost human-AI team accuracy [46, 47, 42, 44, 34, 30, 16, 32] and raises needs for future research. We open-source CHM-Corr++, an interactive tool for editing image classifier attention (see an interactive demo here). We release code and data on github.
@@ 1 Introduction
Despite much attention from the community, the practical utility of Explainable AI (XAI) tools in downstream applications (e.g., image classification [20, 24, 37, 43]) remains limited, hindering human-AI collaboration in real-world settings. A major limitation is that there is no interface for humans to provide feedback to the model so that it can update its decisions, which could change users’ thoughts and final decisions. For example, feature importance maps [9, 15] and example-based explanations [37] are among the most popular XAI methods that offer insights into “what a model is looking at” and which real examples support a model decision, respectively. However, they only offer a static, one-time explanation of the input. Here, we test allowing users to edit the “attention” input to the model and observe updated model decisions iteratively until they are ready to make the final decisions (Fig. 1).
@@
We perform our study on CHM-Corr, a recent explainable CUB-200 bird classifier [48] that combines the best of both worlds by first finding, at the patch level, how the input image is similar to the nearest training-set examples, and then using these patch-wise correspondences to predict the image label (see Fig. 2b and Supp. A3 for examples of CHM-Corr explanations). CHM-Corr explanations enabled users to achieve state-of-the-art human-AI team accuracy in bird identification on CUB-200 [48].
We build an interactive interface called CHM-Corr++ for CHM-Corr, allowing users to manipulate the “attention” of the CHM-Corr classifier by selecting the set of patches that the classifier uses in its decision-making step (Fig. 1). By iteratively telling the model where to look, and observing the changes in the output space (see Fig. 2), users could better understand the AI model and make more informed and accurate decisions.